# PTEB: Towards Robust Text Embedding Evaluation via Stochastic Paraphrasing at Evaluation Time with LLMs

### v1

@@ 5 Conclusion
We introduce PTEB, a sentence embedding evaluation method that stochastically paraphrases MTEB data at eval-time. We show that the paraphrase LLMs yield semantically preserving yet token-diverse outputs. Since our end-to-end LLM-based method indirectly grounds the LLMs in human gold ratings, it can serve as a template for researchers who cannot include expensive human evaluations. Building on this, we demonstrate that the average performance of the tested embedding models drops across 7 MTEB tasks when paraphrased, demonstrating that PTEB effectively uncovers generalisation failures. As a multilingual check, we included 3 multilingual datasets covering 10 languages and 3 different tasks. Ablations over LLM judges and embedding models with different parameter counts show clear runtime–quality trade-offs for the models in scope, and that smaller models (generative LLMs: 12B; sentence encoders: 4B) perform comparably to larger ones on our tests. Our experiments confirm that PTEB yields statistically robust results across multiple runs - even if the paraphrase prompt is varied. In general, PTEB is a novel and more robust benchmarking methodology well suited to augment current embedding benchmarks.
@@ Acknowledgments
PTEB builds upon MTEB (Muennighoff et al., 2023) and MMTEB (Enevoldsen et al., 2025), released under the Apache License 2.0. PTEB is independent work and is not affiliated with or endorsed by the original MTEB or MMTEB authors.
We thank the MTEB maintainers and contributors for the benchmark suite and for making the code publicly available as well as the developers from Hugging Face and of Sentence Transformers for enabling open research.
We also thank the developers of the open-weight models used in this work, including Gemma-3 (Gemma Team et al., 2025; Vera et al., 2025), Qwen3 (Yang et al., 2025), and GPT-OSS (OpenAI et al., 2025).
We acknowledge the use of AI-assisted coding (ChatGPT with GPT-5; Claude Code with Sonnet 4.0/4.5 and Opus 4.1).
6 Limitations
The main limitation of our work lies in the restriction of non-STS datasets to one per task and a single evaluation run. This choice was deliberate: our objective was to demonstrate the applicability of our approach and to provide initial insights rather than to present a comprehensive benchmark of embedding models like MTEB does. Moreover, our analyses of STS datasets using $n=6$ runs already demonstrated the statistical robustness of PTEB.
For the same reason, we limited our evaluation on PTEB to five embedding models and 3 multilingual datasets. In future work, we are planning to implement PTEB as an actual benchmark and include more embedding models and datasets. This is not to replace MTEB but to provide additional signal for embedding model performance.
Furthermore, we did not conduct human evaluations to validate the generated paraphrases. Our goal in this methodological work is an end-to-end LLM-based protocol in which paraphrase models are judged by LLMs calibrated to STS gold ratings, providing an indirect human anchor. This design democratises evaluation by enabling researchers and practitioners with limited budgets to run robust studies without costly, labour-intensive human annotation, while acknowledging that targeted human audits remain valuable for high-stakes settings. Nevertheless, going forward, we plan to include human evaluations to further validate our LLM judge-based approach.
Moreover, the evaluations of the LLM judge and paraphrase models were limited to a single run. This was sufficient for our goal of demonstrating an end-to-end methodology for PTEB that grounds generative model selection in scientific rigor. For production scenarios, our method can serve as a template and be expanded accordingly. Alternatively, one might pragmatically select a paraphrase model directly.
7 Ethical Considerations
We did not identify major ethical concerns arising from this work. However, LLMs (and, hence, PTEB) may replicate or amplify biases present in their training data (Bender et al., 2021). To support ethical scrutiny and reproducibility, we use only open-weight generative models and only embedding models available via Sentence Transformers (i.e., no commercial APIs), enabling researchers to reproduce results, pin versions, and audit behaviour including limited inspection of internals without API accessibility limitations or usage restrictions. While it does not by itself resolve bias, it improves assessability for ethical research.
No personal data was collected and all evaluations are based on public datasets.
Appendix A Non-STS Datasets and Metrics
The following English non-STS datasets (evaluation metrics) were used:
  • Classification: Banking77 (Casanueva et al., 2020) and AmazonCounterFactuals (O’Neill et al., 2021) (Accuracy)
  • Clustering: TwentyNewsGroups Lang (1995) (V-measure)
  • Pair Classification: TwitterSemEval (Xu et al., 2015) and RTE3 Giampiccolo et al. (2007) (average precision)
  • Reranking: AskUbuntuDupQuestions (Wang et al., 2021a) (MAP)
  • Retrieval: ArguAna (Boteva et al., 2016) (nDCG@10)
  • Summarisation: SummEval Fabbri et al. (2021) (Spearman correlation)
@@ Appendix A Non-STS Datasets and Metrics
Further details on all datasets are available in Muennighoff et al. (2023) and Enevoldsen et al. (2025), as well as on https://huggingface.co/mteb/datasets.
Appendix B Implementation Details
@@
Also, for task-specific hyperparameters we applied the MTEB-defaults (v1.38.33), e.g. max_iter=100 for the Logistic Regression model in the Classification task. For a description of the evaluation protocol for each task the reader may refer to Muennighoff et al. (2023). MTEB is licensed under the Apache License 2.0.
  • qwen3:32b
    • Quantisation: Q4_K_M
    • License: Apache License Version 2.0, January 2004
  • gpt-oss:20b
    • Quantisation: MXFP4
    • License: Apache License Version 2.0, January 2004
  • gemma3:27b
    • Quantisation: Q4_K_M
    • License: Gemma Terms of Use (last modified 2024-02-21)
We used the default hyperparameters and environment variables (e.g. OLLAMA_FLASH_ATTENTION=1). Further information on the models is available at www.ollama.com/models.
@@ C.1 LLM Judge Prompts
C.2 Paraphrase Evaluation Prompts
@@ C.3 Paraphrase Prompts
Used in step 3 of our method to test sensitivity to prompt changes. The variants were obtained by asking GPT-5 to rephrase the default prompt:
@@
Appendix D Non-English Languages
The following language pairs are included in the multilingual datasets. Since AmazonCounterFactuals is not a paired dataset, only a single language is given. All languages are encoded in Alpha-2 Country Codes. [^6] https://www.iban.com/country-codes
AmazonCounter-
Factuals
RTE3 STS17
EN DE–DE EN–EN
DE EN–EN AR–AR
JP FR–FR EN–AR
IT–IT EN–DE
EN–TR
ES–EN
ES–ES
FR–EN
IT–EN
KR–KR
NL–EN
Table 11: Languages in AmazonCounterfactuals, RTE3, and STS17.
Appendix E Additional Analyses

### v2

@@ 5 Conclusion
We introduce PTEB, a sentence embedding benchmark that stochastically paraphrases MTEB data at evaluation time. Using LLM judges and human evaluation, we show that the LLMs in scope yield semantics-preserving yet token-diverse paraphrases. Building on this, we demonstrate that the average performance of the tested embedding models drops across 7 NLP tasks when paraphrased, showing that PTEB effectively uncovers generalisation failures. For non-English languages, we included 6 multilingual datasets covering 25 languages incl. low resource (e.g., African) languages. Ablations over LLM judge and sentence embedding model sizes show runtime–quality trade-offs for the models in scope and that smaller models (12B parameters for LLM judges and 4B for embedding models) perform comparably to larger ones in our study. Our experiments confirm that PTEB yields statistically robust results across multiple runs with different random seeds and even if the paraphrase prompt is varied. In general, PTEB is a novel, more robust evaluation methodology well suited to augment current embedding benchmarks.
6 Limitations
LLM judge and paraphrase model evaluation. The evaluations of the LLM judge and paraphrase models were limited to a single run. This was sufficient for our goal of demonstrating an end-to-end methodology for PTEB that grounds generative model selection in human gold ratings without human evaluation. For production scenarios, our method can serve as a template and be expanded accordingly. Alternatively, one might pragmatically select a paraphrase model directly.
Human validation. The human evaluation in our work is limited to 8 languages and 400 sentence pairs (each rated by 2 evaluators). For paraphrasing of English datasets we relied on an automated approach using LLM judges which were only indirectly validated by grounding them in STS gold labels but not validated through human evaluation.
PTEB paraphrase model. Our embedding model evaluation relies exclusively on gemma3-27b paraphrases. While the cross-model-family validation (Table 5) confirmed consistent semantic similarity scores across different LLM judges, future work should examine whether the paraphrasing style itself affects embedding model rankings by running PTEB with different paraphrasers.
In an initial analysis using the STSB dataset, we compared the edit distance between paraphrases and original texts to the performance drops on PTEB for gemma3-27b, gpt-oss-20b, and qwen3-32b (see Figure 8 in Appendix F). We found that generative models that generate paraphrases with larger edit distances are associated with larger performance drops of embedding models on PTEB. That is, in our initial analysis 4 out of 5 embedding models show the largest performance drops when using gemma3-27b as a paraphrase model (which is the paraphraser generating the largest EDs; see Table 4). Similarly, qwen3-32b, which produces the lowest EDs, yields the smallest performance drops for 4 out of 5 embedding models. Despite the need for further analysis, we interpret this as initial evidence that using the ED as a selection criterion is essential, since our goal was to generate paraphrases with substantial token-level changes. Which was exactly one of the reasons to choose gemma3-27b (see Table 4) as also illustrated in a non-cherry picked example in Figure 5.
7 Ethical Considerations
Model bias. However, LLMs (and, hence, PTEB) may replicate or amplify biases present in their training data (Bender et al., 2021). To support ethical scrutiny and reproducibility, we use only open-weight models (i.e., no commercial APIs), enabling researchers to reproduce results, pin versions, and audit behaviour including limited inspection of internals without API accessibility limitations or usage restrictions. While it does not by itself resolve bias, it improves assessability for ethical research.
Personal data. No personal data was collected and all evaluations are based on public datasets.
Crowd worker compensation. We ensured appropriate and equitable compensation by paying all annotators an appropriate and identical rate regardless of their geographic location. At \$0.25 per sentence pair, it corresponds to an estimated hourly wage of \$36.
Overall, we did not identify major ethical concerns arising from this work.
@@ Acknowledgments
We thank the authors of the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023) and the Massive Multilingual Text Embedding Benchmark (MMTEB) (Enevoldsen et al., 2025) for making their code and the included datasets publicly available. While PTEB builds on MTEB and MMTEB, it is an independent extension and is not affiliated with or endorsed by the MTEB or MMTEB authors.
We also thank the anonymous reviewers and human evaluators for their valuable contributions.
We acknowledge the use of AI, such as Anthropic’s Claude Code, OpenAI’s ChatGPT, and Google’s Gemini for assisted coding and writing, e.g., for improving the language of our paper.
This research was partially supported by the Horizon Europe project GenDAI (Grant Agreement ID: 101182801) and by the ADAPT Research Centre at Munster Technological University. ADAPT is funded by Taighde Éireann – Research Ireland through the Research Centres Programme and co-funded under the European Regional Development Fund (ERDF) via Grant 13/RC/2106_P2.
Appendix A Datasets
@@
Table 13 shows all PTEB datasets with the respective metrics and the languages included in this study. Further details on all datasets are available in Muennighoff et al. (2023) and Enevoldsen et al. (2025), as well as at https://huggingface.co/mteb/datasets.
Tasks/Datasets Languages Metric
Classification Accuracy
Banking77 (Casanueva et al., 2020) eng
AmazonCounterfactuals (O’Neill et al., 2021) deu, eng, jpn
Clustering V-measure
TwentyNewsGroups (Lang, 1995) eng
MasakhaNEWSClusteringS2S (Adelani et al., 2024) amh, eng, fra, hau, ibo, lin, lug, orm, pcm, run, sna, som, swa, tir, xho, yor
Pair Classification Average Precision
TwitterSemEval2015 (Xu et al., 2015) eng
RTE3 (Giampiccolo et al., 2007) deu, eng, fra, ita
Reranking MAP
AskUbuntuDupQuestions (Wang et al., 2021a) eng
RuBQ Reranking (Rybin et al., 2021) rus
Retrieval (Due to corpus length only queries are paraphrased.) nDCG@10
ArguAna (Boteva et al., 2016) eng
TwitterHjerne (Holm, 2024) dan
Summarisation Spearman corr.
SummEval (Fabbri et al., 2021) eng
Semantic Textual Similarity (STS) Spearman corr.
BIOSSES (Soğancıoğlu et al., 2017) eng
SICK-R (Marelli et al., 2014) eng
STS12 (Agirre et al., 2012) eng
STS13 (Agirre et al., 2013) eng
STS14 (Agirre et al., 2014) eng
STS15 (Agirre et al., 2015) eng
STS17 (Cer et al., 2017) eng, ara, deu, fra, ita, nld, spa, tur
STS22 (Chen et al., 2022) eng
STSBenchmark (Cer et al., 2017) eng
Table 13: Overview of datasets included in PTEB. For each dataset, we report the language(s) and evaluation metric; since we only used the English subset of STS22, other languages are not listed for this dataset.
Appendix B Human Evaluation Details
For each language, we specified the most relevant countries and filtered the Amazon Mechanical Turk workers based on their location. Each rating/HIT was paid \$0.25. Based on our assumption that a rating/HIT takes on average 25 seconds, this equals a payment of \$36 per hour. This payment was independent of the worker’s location. The time limit for each worker to complete a batch was set to 1 hour. For each language we ran an initial pilot with 20 sentence pairs. In the case of one language, Danish, we were not able to collect any ratings and therefore replaced it with French - while not being a low-resource language, French is a highly relevant for PTEB since 3 of the datasets in this study include French text (see Table 13).
As a basic validation of the human ratings, we included dummy sentence pairs using an almost identical sentence pair (easy positive) and a completely dissimilar sentence pair (easy negative) in each batch. For each worker, we checked if these were rated correctly within a margin of ±1 before accepting their ratings. Overall, we employed the following validation steps for all ratings based on our 0–5 rating scale:
  1. 1.
    Check the ratings of the two dummy sentences. [^9] Note that not all workers completed the dummy sentence pair ratings.
  2. 2.
    Prompt GPT-5.1 and Claude Sonnet 4.5 with the sentence pairs and human ratings and ask both to flag any ratings which deviate by more than 2.0 from their own similarity estimate.
  3. 3.
    Identify any sentence pair with a similarity rating of 2.0 and below.
  4. 4.
    Translate all sentence pairs identified in step 1 to 3 to English using Google Translate and manually inspect the validity of the ratings.
  5. 5.
    Reject workers that fail the dummy sentence pairs within the pre-defined acceptability threshold or repeatedly submitted ratings that deviate by more than 2 from the similarity rating based on the manual inspection and the LLM ratings. Validate the potential candidates to be dropped using the results from step 4.
These validation steps and criteria have been defined a priori to avoid the common mistake of ad-hoc rejection of workers or ratings leading to biased results (Thomson et al., 2024). Generally, we aimed to keep rejections to a minimum.
Based on these steps, the ratings of exactly one worker for Arabic were rejected since they incorrectly rated the dummy sentence pairs beyond the acceptability threshold. Also, the ratings of the worker in question were consistently more than 2 points below that of other Arabic workers, the two LLMs estimates, and the manual translation-based approach.
Furthermore, we obtained clarification by two additional workers whose ratings were flagged during the above process while they correctly rated the dummy sentences. Following this, their ratings were ultimately approved.
Appendix C Implementation Details
We used the default hyperparameters and environment variables (OLLAMA_FLASH_ATTENTION=1, for example). Further information on the models is available at www.ollama.com/models.
@@ Appendix C Implementation Details
For retrieval datasets only the queries were paraphrased due to the length of the corpora. Also, for task-specific hyperparameters we applied the MTEB-defaults (v1.38.33), e.g. max_iter=100 for the Logistic Regression model in the Classification task. For a description of the evaluation protocol for each task the reader may refer to Muennighoff et al. (2023). MTEB is licensed under the Apache License 2.0.
@@ D.1 LLM Judge Prompts
D.2 Paraphrase Evaluation Prompts
Used in step 2 of our method to rate the paraphrase models:
@@ D.3 Paraphrase Prompts
D.4 Paraphrase Prompt Variatons for Sensitivity Analysis
Used in the prompt sensitivity analysis to test sensitivity to prompt changes. The variants were obtained by promtping GPT-5 to rephrase the default prompt:
@@ D.4 Paraphrase Prompt Variatons for Sensitivity Analysis
Appendix E Run-Aggregated Statistical Analysis
Table 14 shows the results of the statistical analysis when taking the average over the $n=6$ runs (which reduces the number of observations from $9\times 6=54$ to just $9$) before calculating the HL estimated location shift and applying the Wilcoxon signed-rank test. Due to the smaller number of observations we do not calculate bootstrap CIs.
all-mpnet-base-v2 vs. $\Delta_{HL}$ $p_{Holm}$
embeddinggemma-300m +7.17 0.02
mxbai-embed-large-v1 -3.62 0.04
e5-mistral-7b-instruct -3.42 0.04
qwen3-embedding-8b -4.46 0.02
Table 14: Run-aggregated comparison of encoders on PTEB STS vs. all-mpnet-base-v2 as baseline. $\Delta_{HL}$ is the Hodges–Lehmann estimate of the location shift with its 95% confidence interval; negative values favour the listed model. P-values are exact Wilcoxon with Holm correction. (9 datasets, average over $n=6$ runs first)
The run-aggregation follows Søgaard et al. (2014), who average in their experiments for 2 different NLP tasks over 3 and 10 runs respectively.
Recommended minimum sample sizes for the bootstrap method vary considerably, from 200 (Søgaard et al., 2014) to 20 (Colas et al., 2018), with the requirement that samples be representative of the population (Colas et al., 2018; Dror et al., 2018). In our experiments, the number of observations is $9\times 6=54$ when pooling over datasets and runs (see Table 7) and $9$ when averaging across runs first (see Table 14). Given these modest number of observations, we use the Wilcoxon signed-rank test for significance testing.
The results show that even in this conservative analysis, the p-values remain below $\alpha=0.05$ after Holm-correction.
Appendix F Additional Analyses
Figure 7 visualises how the performance of embeddings models drops on PTEB vs. original data per for the various task.
@@ Appendix F Additional Analyses
Figure 8 shows how paraphrase models that generate paraphrases with larger edit distance compared to the original data (see Table 4) tend to lead to larger performance drops for embedding models on the STSBenchmark dataset.
Figure 8: Edit distance of the paraphrases generated by the different generative paraphrase models (X-axis) vs. the performance drop of the embedding models on PTEB compared to original data for STSBenchmark ($n=1$ runs; in %).